# Undesirable Memorization in Large Language Models: A Survey

### v1

@@ Abstract
While recent research increasingly showcases the remarkable capabilities of Large Language Models (LLMs), it’s vital to confront their hidden pitfalls. Among these challenges, the issue of memorization stands out, posing significant ethical and legal risks. In this paper, we presents a Systematization of Knowledge (SoK) on the topic of memorization in LLMs. Memorization is the effect that a model tends to store and reproduce phrases or passages from the training data and has been shown to be the fundamental issue to various privacy and security attacks against LLMs.
We begin by providing an overview of the literature on the memorization, exploring it across five key dimensions: intentionality, degree, retrievability, abstraction, and transparency. Next, we discuss the metrics and methods used to measure memorization, followed by an analysis of the factors that contribute to memorization phenomenon. We then examine how memorization manifests itself in specific model architectures and explore strategies for mitigating these effects. We conclude our overview by identifying potential research topics for the near future: to develop methods for balancing performance and privacy in LLMs, and the analysis of memorization in specific contexts, including conversational agents, retrieval-augmented generation, multilingual language models, and diffusion language models.
Index Terms:
Memorization, Large Language Models, Privacy in LLMs
@@ I Introduction
In recent years, large language models (LLMs) have demonstrated remarkable advancements, driven by the scaling of model parameters, large amounts of data, or training paradigms [1, 2, 3, 4]. State-of-the-art models have exhibited capabilities across a broad spectrum of Natural Language Processing (NLP) tasks, consistently pushing the benchmark records in areas such as text generation, code synthesis, machine translation, question answering, and summarization [5, 6]. These models use self-supervised learning on massive datasets, enabling them to perform competitively or even surpass human-level performance in specific tasks [7, 8].
@@
Despite these impressive advancements of LLMs, researchers have pointed out the problematic features of these models, including hallucination [9], bias [10], and privacy and security [11]. In the context of data privacy, memorization is one of the most significant issues. Memorization in LLM refers to the model’s tendency to store and reproduce exact phrases or passages from the training data rather than generating novel or generalized outputs. While memorization can be advantageous in knowledge-intensive benchmarks, such as factual recall tasks or domain-specific question answering [12], it introduces significant ethical and legal challenges: First, models may inadvertently reveal sensitive or private information included in their training data, posing significant privacy and security risks [13, 14]. Second, the ability of LLMs to repeat verbatim copyrighted or proprietary text from their training data raises issues related to intellectual property infringement [15, 16].
@@
These challenges motivate the need to further explore memorization in LLMs to effectively tackle its associated challenges. We provide an overview of aspects related to memorization, emphasizing the need for further exploration of this topic, thereby balancing model performance with the risks of privacy breaches and ethical concerns.
@@
Before recent advances in LLMs, memorization was explored extensively as a topic in machine learning and deep learning. Usynin et al. [17] explore memorization in machine learning. They propose a framework to quantify the influence of individual data samples and detect memorization in various learning settings. Wei et al. [18] provide a systematic framework for understanding memorization in deep neural networks, discussing LLM memorization as a type of deep neural network. Survey papers on the privacy and safety of LLMs often address memorization as a core phenomenon, framing it as both a privacy issue and a foundational factor that supports other challenges they explore [19, 20]. Hartmann et al. [21] provide an overview of memorization in general-purpose LLMs. They aggregate memorization-related topics from the copyright, privacy, security, and model performance perspectives. In our survey, we not only add more recent work, but we specifically focus on memorization as an undesirable phenomenon, with the aspects intentionality, severity of memorization, retrievability, and transparency of memorization. In addition, we provide an extensive and concrete research agenda for the near future. Table I provides a comparison of our scope to the scopes of previous surveys.
@@
Usynin et al. [17] Memorization in general machine learning We focus on LLMs
@@
Neel and Chang [19] Privacy Problems in LLMs Memorization is treated as an instance of a privacy related problems rather than being the focus.
@@
Smith et al. [20] Privacy Problems in LLMs Memorization is treated as an instance of a privacy related problems rather than being the focus.
@@
Hartmann et al. [21] Memorization in general-purpose LLMs We focus on unintended and undesirable memorization in LLMs, and propose concrete directions for future work.
TABLE I: List of existing surveys vs our work
@@
In the following sections of this paper, we provide a comprehensive exploration of memorization in LLMs, structured into six main sections. We begin with the Spectrum of Memorization[II], where we examine the concept from multiple perspectives including intentionality, retrievability, severity, abstraction, and transparency, offering a deep understanding of how memorization occurs. Next, in Measuring Memorization [III], we review various methodologies that have been used in previous studies to quantify and assess memorization within LLMs. We then explore the Influences and Dynamics of Memorization [III], identifying key factors and conditions that contribute to this phenomenon. Following this, in section [V] we analyze how memorization appears in specific LLM architectures. Next, in the Mitigating Memorization section [VI], we discuss strategies and techniques employed to minimize or control memorization in LLMs, addressing concerns related to privacy and generalization. Finally, we conclude with Future Directions and Open Challenges [VII], outlining potential areas for further research and unresolved questions, before summarizing our findings in the Conclusion [VIII]. Figure 1 shows a visual overview of our systemization scope.
Figure 1: Graphic summary of our systematization of knowledge
@@ II Spectrum of Memorization
II-A Intentionality
Although memorization of factual information can be helpful for the model the model perform more accurately on benchmark tasks such as question answering, other data might be retained without any clear purpose, and this might cause issues with privacy and copyright. This raises questions about whether certain instances of memorization is deliberate (by design) or incidental and undesirable. Understanding the difference between deliberate memorization and unintended retention is crucial for both improving model design and addressing issues like privacy and data security [22, 23, 24].
Unintended memorization
Carlini et al. [25] explores the concept of unintended memorization in neural networks, identifying it as a form of memorization that occurs independently of overfitting and often early in the training process, before the learning has fully converged. To understand and measure this phenomenon, they introduced the “exposure metric” as part of a testing methodology (see Section III-A). It is important to note that the language modeling loss function itself is chosen in a way to be optimal when the language model’s output are recreations of training data, however, for the unintended memorization, this happens even when the loss function still has some distance to its optimal value. Unintended memorization could lead to various issues including:
  • Security vulnerabilities: Unintended memorization of confidential information like API keys or passwords could lead to security breaches [25].
  • Copyright violation: Memorization of copyrighted material may lead to legal challenges, especially if the model can reproduce substantial portions of protected works [16].
Deliberate memorization
Deliberate memorization, in contrast, refers to the intentional design of LLMs to retain specific types of information from their training data. This is often a desirable feature, as it allows models to store and utilize vast amounts of knowledge. Key aspects of deliberate memorization include:
  • Knowledge retention: LLMs are often deliberately designed to memorize facts, concepts, and general knowledge to enhance their performance in tasks like question answering and information retrieval [26].
As Ranaldi et al. [28] show in their experiments, deliberate memorization is generally beneficial for model performance, however, it also raises important considerations:
  • Bias and fairness: The choice of what information to deliberately memorize can introduce or amplify biases in the model output.
  • Transparency and auditing: Understanding what has been deliberately memorized is crucial for model interpretability and auditing purposes.
II-B Severity of memorization
While memorization in LLMs is widely discussed, there is no universally accepted definition. The definition of memorization can vary based on factors such as the level of detail required, the context in which the information is recalled, and the specific task or application at hand. This section discusses different severity levels of recall that are explored in the literature evolving around memorization in LLMs, including perfect memorization, verbatim memorization, and approximate memorization.
@@ II-B Severity of memorization
Perfect memorization refers to a scenario where a model retains and can reproduce all training data exactly. Kandpal et al. [29] define Perfect Memorization as follows:
Definition (Perfect Memorization).
A model is said to exhibit perfect memorization behavior if it only assigns non-zero probability to the samples it has been trained on. Sampling from a perfect memorization model is identical to sampling from the training data.
This form of memorization acts as an imaginary language model and is used as a baseline for the experiments.
@@
Verbatim memorization is a more specific form of memorization that involves the exact reproduction of strings from the training data. It captures instances where the model outputs a training sample without any alterations. This type of memorization is straightforward to identify and is often the primary focus in discussions about memorization in language models [30, 13].
Carlini et al. [13] define verbatim memorization under the “Eidetic Memorization” terminology as follows:
Definition (Verbatim memorization).
a string $s$ from the training set is defined as verbatim memorized if there exists a prompt $p$ so that $LM(p)=s$.
Tirumala et al. [31] consider another version of this as “Exact memorization”:
@@
[Exact memorization] a context $c=(p,s)$ from the training set is being memorized if $Argmax(LM(p))=s$.
This definition is more strict and does not capture the memorization under different forms of sampling methods, rather it only considers the greedy sampling [IV-E].
It should be noted that these definitions do not impose any restrictions on the length of prompt $p$ or the generation $s$. However, if the generated text is too short, it may not be appropriate to classify it as memorization. Typically, a certain number of tokens is considered when assessing memorization; for example, [23, 31] use a 50-token output restriction.
@@
Ippolito et al. [32] define “Approximate memorization” as follows:
Definition (Approximate memorization).
@@
a suffix $s$ for prefix $p$ is labeled as approximately memorized if for generation $g=LM(p)$, $BLEU(g,s)>0.75$.
They discuss that this threshold was chosen based on qualitative analyses of the samples. Examples of this are shown in Table 2. By adopting this definition of memorization, they show that the measurement of memorization can increase by a factor of two. However, they also note that this definition can lead to both false positives and false negatives when compared to human judgment, indicating a potential direction for future investigations.
Figure 2: BLEU similarity score of more than 0.75 (from Ippolito et al. [32])
II-C Retrievability
@@
The common way to analyze memorization in LLMs is through sampling. This involves selecting tokens probabilistically from the model’s predicted distribution at each step, to see if a sequence of probabilistically generated tokens can be traced back to the training data. Since the LLM’s input and output domains are both discrete, it is important to note that it might not be possible to extract all of the memorized content through given a finite number generation trials through sampling. Based on how we can make an LLM output the training data we can classify memorization into extractable and discoverable.
@@
Extractable memorization refers to the ability to retrieve specific information from a model’s training data without direct access to that data. Carlini et al. [23] defines “Extractable memorization” as follows:
@@
(Extractable Memorization) Given a model with a generation routine $Gen$, an example $s$ from the training set $S$ is extractably memorized if an adversary (without access to $S$) can construct a prompt $p$ that makes the model produce $s$ (that is, $Gen(p)=s$) .
A more specific form of this is k-extractable memorization:
Definition.
(K-extractable Memorization) A string $s$ is said to be k-extractable if it (a) exists in the training data, and (b) is generated by the language model by prompting with $k$ prior tokens [33].
@@
Discoverable memorization measures the extent to which models can reproduce their training data when explicitly prompted with data from their training set. Nasr et al. [34] suggests the following definition for “Discoverable memorization”:
@@
[…]
Discoverable memorization provides an upper bound for data extraction, compared to extractable memorization’s lower bound. Ideally, discoverable memorization requires querying the model with its entire training set, which is computationally intractable. Also, noteworthy is that discoverable memorization differs from extractable memorization in that the prompt $p$ is known to be from the training set.
@@
Comparing discoverable and extractable Memorization
@@
II-D Abstraction
Abstraction in the context of memorization refers to the level of specificity at which information is retained and reproduced by language models. This spectrum ranges from highly specific, fact-based memorization to more generalized concept-level retention. This is related to the concept of verbatim memorization in Section II-B, but abstraction focuses more on information (semantics of facts) than on literal word use. In terms of abstraction, memorization can be categorized into two main types: factual memorization and conceptual memorization. These categories represent different levels of information granularity and have different implications for model performance, privacy concerns, and potential applications.
Factual memorization
Factual memorization in language models involves accurately retaining and reproducing specific pieces of information from training data, often relying on context cues to recall details such as names, dates, and numbers or the relations between such entities. Petroni et al. [12], Jiang et al. [38], AlKhamissi et al. [39], Luo et al. [40] investigate the measurement of factual knowledge in LLMs by evaluating their comprehension of facts from knowledge bases using ranking metrics. Pezeshkpour [41] proposes two approaches of explicit and implicit instillation of factual knowledge in LLMs while proposing to use KL-divergence as a metric for measuring factual memory. Yu et al. [42] show that the frequency of facts in the pretraining data significantly influences a language model’s behavior. They find that models are more likely to generate memorized capital cities for frequently mentioned countries, while being less responsive to in-context information. Additionally, they demonstrate that larger models tend to favor memorized answers over in-context details, even when the memorized facts are less frequent. Kang and Choi [43] also show that factual knowledge probing accuracy in large language models is linked to subject-object co-occurrence. They reveal that LLMs often prioritize frequently co-occurring word pairs, which can override correct answers, particularly when those answers are infrequent. Their findings highlight the need to address co-occurrence bias in future model development.
While factual memorization could be beneficial for the performance of LLMs in tasks namely question answering, it could potentially lead to various issues such as undesired bias and privacy and copyright violations.
Conceptual memorization
Conceptual memorization involves the model’s ability to internalize and generalize broader ideas, patterns, and concepts from its training data [44, 45, 46, 4]. That is, Lee et al. [47] explore the idea plagiarism of GPT-2 model and empirically show that language models go beyond merely generating verbatim text from their training data; rather they also rephrase sentences and imitate ideas from different sources.
Compared to factual memorization, conceptual memorization generally poses a lower risk of exposing specific, sensitive information. At the same time, it is more difficult to be detected and it may lead to perpetuation of biases present in the training data at a more systemic level.
The boundary between conceptual memorization and the understanding of the model (generalization) is subtle and still debated.

### v2

@@ Abstract
While recent research increasingly showcases the remarkable capabilities of Large Language Models (LLMs), it is equally crucial to examine their associated risks. Among these, privacy and security vulnerabilities are particularly concerning, posing significant ethical and legal challenges. At the heart of these vulnerabilities stands memorization, which refers to a model’s tendency to store and reproduce phrases from its training data. This phenomenon has been shown to be a fundamental source to various privacy and security attacks against LLMs.
In this paper, we provide a taxonomy of the literature on LLM memorization, exploring it across three dimensions: granularity, retrievability, and desirability. Next, we discuss the metrics and methods used to quantify memorization, followed by an analysis of the causes and factors that contribute to memorization phenomenon. We then explore strategies that are used so far to mitigate the undesirable aspects of this phenomenon. We conclude our survey by identifying potential research topics for the near future, including methods to balance privacy and performance, and the analysis of memorization in specific LLM contexts such as conversational agents, retrieval-augmented generation, and diffusion language models.
Given the rapid research pace in this field, we also maintain a dedicated repository of the references discussed in this survey [^1] https://github.com/alistvt/undesirable-llm-memorization, which will be regularly updated to reflect the latest developments.
@@ I Introduction
In recent years, large language models (LLMs) have demonstrated remarkable advancements, driven by the scaling of model parameters, large amounts of data, and extensive training paradigms [1, 2, 3, 4]. State-of-the-art models have exhibited capabilities across a broad spectrum of natural language processing (NLP) tasks, consistently pushing the envelope in areas such as text generation, code synthesis, machine translation, question answering, and summarization [5, 6]. These models are trained on massive datasets, enabling them to perform competitively or even surpass human-level performance in specific tasks [7, 8].
@@
Despite these impressive advancements of LLMs, researchers have shown that there are certain problems with these models, including hallucination [9], bias [10], and privacy and security vulnerabilities [11]. In the context of data privacy, memorization is one of the core sources of concern. Memorization in LLM refers to the model’s tendency to store and reproduce exact phrases or passages from the training data rather than generating novel or generalized outputs. While memorization can be advantageous in knowledge-intensive benchmarks, such as factual recall tasks or domain-specific question answering [12], it also introduces ethical and legal challenges: models may inadvertently reveal sensitive or private information included in their training data, posing significant privacy and security risks [13, 14]. In addition, the ability of LLMs to repeat verbatim copyrighted or proprietary text from their training data raises issues related to intellectual property infringement [15, 16].
@@
These challenges motivate the need to further explore memorization in LLMs to effectively tackle the associated challenges. In this paper, we provide an overview of aspects related to memorization, emphasizing the need for further exploration of this topic, thereby balancing model performance with the risks of privacy breaches and ethical concerns.
I-A Related surveys
@@
Before recent advances in LLMs, memorization has been explored extensively as a topic in machine learning and deep learning mostly with a focus on security and privacy. Usynin et al. [17] explore memorization in machine learning. They propose a framework to quantify the influence of individual data samples and detect memorization in various learning settings. Wei et al. [18] provides a systematic framework for understanding memorization in deep neural networks, discussing LLM memorization from the view of deep neural networks. Survey papers on the privacy and safety of LLMs often address memorization as a core phenomenon, framing it as both a privacy issue and a foundational factor that supports other non-security/privacy-related challenges they explore [19, 20]. Hartmann et al. [21] provide an overview of memorization in general-purpose LLMs. They aggregate memorization-related topics from the copyright, privacy, security, and model performance perspectives. In our survey, we not only incorporate more recent work but also specifically focus on memorization as an undesirable phenomenon, examining it with the aspects of granularity, retrievability, and desirability. Our work is not only relevant for privacy but also for (un)safety and bias as undesirable properties of LLMs when memorization takes place. In addition, we provide an extensive and concrete research agenda for the near future. Table I provides a comparison of the scope of our survey with that of previous surveys.
@@
Usynin et al. [17] Memorization in machine learning We focus on LLMs
@@
Neel and Chang [19] Privacy Problems in LLMs Memorization is treated as one of the sources of privacy-related problems rather than being the focus.
@@
Smith et al. [20] Privacy Problems in LLMs Memorization is treated as one of the sources of privacy-related problems rather than being the focus.
@@
Hartmann et al. [21] Memorization in general-purpose LLMs We focus on undesirable memorization in LLMs, and propose concrete directions for future work.
TABLE I: List of existing surveys vs. our work
I-B Data selection
The selection of works included in this survey was guided by the goal of capturing a comprehensive view of the literature surrounding memorization in LLMs, with a particular focus on its undesirable aspects. However, to provide a well-rounded taxonomy and to contextualize these challenges, we also consider works that explore the desirable aspects of memorization, such as generalization and knowledge retention, as they offer complementary insights into the mechanisms at play.
Selection process
The process of identifying relevant papers was iterative, beginning with the most widely cited early studies by recognized researchers in the field (14 papers). These foundational and influential works were selected as they represent key milestones that have shaped current understandings of memorization in LLMs. Then we included the papers drawn from keyword searches in article repositories. We used Google Scholar [^2] https://https://scholar.google.com/ and arXiv [^3] https://arxiv.org/, as they are predominantly used by researchers in the computer science and computational linguistics domains [22]. For Google Scholar we restricted our keyword search to the papers that include the term “Memorization” in their title and include “Language Model” in their body (memorization is also used as a term in biology). For the scanning of arXiv, we collected the papers that included both of the words “Memorization” and “Privacy” in their abstract. For inclusion, we selected articles published before January 2025.
Additionally, given the broad scope of the topic, our survey draws from adjacent areas, including studies on data extraction, membership inference, and other forms of data leakage, which intersect with the broader concept of memorization. These works were identified through reference chaining, utilizing bibliographies of key papers, and consulting sections of existing survey articles. We believe this approach allows the survey to address the complexities and nuances of memorization in LLMs.
Rejection criteria
After obtaining the initial repository of papers, we manually iterated through the collection and removed non-relevant papers (e.g., out of scope for “undesirable memorization”, focus on deep learning) and the papers that were neither published at a scientific venue nor had any citations. This process and the obtained statistics are summarized in Figure 1.
Figure 1: Our data selection process: Starting with 14 foundational papers as the seed, our literature search expanded through keyword search in Google Scholar (223 papers) and arXiv (111 papers), followed by manual filtering, reference chaining, and exploration of adjacent areas, concluding in a final selection of 99 papers.
I-C Paper organization
This survey is organized to capture the body of literature around undesirable memorization and offers an extensive view of this phenomenon from different perspectives. Our contribution is driven and structured around the following research questions:
  1. RQ1.
    How is memorization defined in the context of LLMs?
  2. RQ2.
    What are the methods used to measure memorization of LLMs?
  3. RQ3.
    What are the factors contributing to memorization?
  4. RQ4.
    What methods are used to prevent or mitigate undesirable memorization?
  5. RQ5.
    What are the important aspects of memorization that are still unexplored or require more research?
@@ I-C Paper organization
In the following sections of this paper, we provide a comprehensive exploration of memorization in LLMs around the mentioned research questions, structured into five main sections. We begin with the Spectrum of memorization (section II), where we examine the concept from multiple perspectives including desirability, retrievability, and granularity, offering a deep understanding of how memorization occurs. Next, in Measuring memorization (section III), we review various methodologies that have been used in previous studies to quantify and assess memorization within LLMs. We then explore the Influencing factors and dynamics of memorization (section III), identifying key factors and conditions that contribute to this phenomenon. Next, in the Mitigating memorization section (section V), we discuss strategies and techniques employed to minimize or control memorization in LLMs, addressing concerns related to privacy and generalization. Finally, we conclude with Future directions and open challenges (section VI), outlining potential areas for further research and unresolved questions, before summarizing our findings in the Conclusion (section VII). Figure 2 shows a visual overview of our survey scope.
Figure 2: Graphic summary of our survey
@@ II Spectrum of Memorization
In this section, we provide an overview of the available definitions for memorization in various dimensions and categorize the existing work according to these dimensions. The overall summary of our findings is presented in Table III along with the other aspects.
II-A Granularity of memorization
While memorization in LLMs is widely discussed in the literature, there is no universally accepted definition for it. The definition of memorization can vary based on factors such as the level of detail required, the context in which the information is recalled, and the specific task or application at hand. This section discusses different levels of granularity of recall that are explored in the literature evolving around memorization in LLMs.
@@ II-A Granularity of memorization
Perfect memorization refers to a setting where a model can generate only from the training data. Kandpal et al. [23] define perfect memorization as follows:
Definition.
[Perfect memorization] A model is said to perfectly memorize its training data if the generation frequencies of sequences are the same as their appearances in the training data. Sampling outputs from a perfect memorization model is identical to sampling from the training data.
Perfect memorization could be viewed as an upper bound for the level of memorization a language model can exhibit. It acts as an imaginary language model to help compare the extent of memorization in practical scenarios.
@@
Verbatim memorization is defined as a form of memorization that involves the exact reproduction of strings from the training data. It captures instances where the model outputs a training sample without any alterations. This type of memorization is straightforward to identify and is often the primary focus in discussions about memorization in language models [24, 13] (Table III).
Carlini et al. [13] define verbatim memorization under the term eidetic memorization as follows:
@@
[Verbatim memorization] Let $LM(p)$ denote the output of a language model when prompted with $p$. A string $s$ from the training set is defined as verbatim memorized if there exists a prompt $p$ such that $LM(p)=s$.
Due to its intuitiveness, various variants of verbatim memorization has been put forward. Tirumala et al. [25] define another version of verbatim memorization as Exact memorization with a focus on the context and sampling method:
Definition.
[Exact memorization] Let $c=(p,s)$ be a context where $p$ is incomplete block that is completed by the sequence $s$. The context $c$ from the training set is considered to be exactly memorized by the language model $LM$ if $Argmax(LM(p))=s$.
This definition is more strict and does not capture the memorization under different forms of sampling methods. It rather considers only the greedy sampling [see Section IV-E].
It should be noted that these definitions do not impose any restrictions on the length of the prompt $p$ or the generated output $s$. However, if the generated text is too short, it may not be appropriate to classify it as memorization. Typically, a certain number of tokens is considered when assessing memorization; 50 tokens length prefix and suffix is used predominantly by the researchers (different employed prompt lengths and generation lengths are summarized in table III).
Throughout the rest of this paper, we use the word memorization interchangeably with the verbatim memorization, as it is the one that is mostly associated with the undesirable aspects of memorization.444discussed in section II-C
@@
Ippolito et al. [26] define Approximate memorization when the generation BLEU similarity score 555a metric used to compare the similarity of texts based on n-grams overlaps [27] with respect to the training data surpasses the 0.75 threshold, which was chosen based on qualitative analyses of the samples. By adopting this definition of memorization, they show that the measurement of memorization can increase by a factor of two compared to only considering verbatim memorization. Similarly, Duan et al. [28] use token-wise Levenshtein as an approximation distance function to experiment with the dynamics of memorized sequences.
Definition.
@@
[Approximate memorization] A suffix $s$ for prefix $p$ is labeled as approximately memorized if for generation $g=LM(p)$, $Sim(g,s)>\delta$; where $Sim(,)$ is a textual similarity metric.
As discussed by Ippolito et al. [26], this definition can lead to both false positives and false negatives when compared to human judgment, indicating a potential direction for future investigations. Since detecting approximate memorization could be resource intensive, Peng et al. [29] introduce a Mini-hash algorithm to efficiently detect approximately memorized content by the LLMs based on the Jaccard similarity metric.
Entity-level memorization
When considering privacy, the relationships between entities and their connections are often more critical than their exact phrasing in a sentence. This means that even approximate memorization might fail to capture certain privacy risks. For example, suppose the sentence “John Smith’s phone number is 012345678.” was included in the training set. If an adversary manages to prompt the language model to generate the following sentence: “I called John Smith yesterday and entered this number into my phone: 012345678,” the model would still be revealing sensitive information. However, this case would not be detected by the memorization granularities discussed earlier. Despite the lack of verbatim memorization, the model is still violating privacy by exposing the association between John Smith and their phone number. Zhou et al. [30] define the entity-level memorization as a phenomenon when the model recalls (generates) an entity when prompted with several other entities, when all of these entities have been linked to each other in the training phase.
Definition.
[Entity memorization] Let $s$ be a training sample containing a set of entities $M$. A prompt $p$ can be constructed to include a strict subset of the entities $N\subset M$. Model is said to show entity memorization if, when prompted with $p$, outputs a response containing some entities in MNM-N.
Zhou et al. [30] conduct various experiments and show that entity-level memorization happens in different scales affected by model size, context length, and repetitions of the entities in the pretraining data. Similarly, Kim et al. [31] target the personally identifiable information (PII) leakage in the OPT models [32] and show that different types of PII can be extracted if the attackers feed intelligently crafted prompts to the model. They also show that the susceptibility of this leakage is influenced by the type of PII; emails and phone numbers show higher extraction rates compared to addresses.
Content memorization
Beyond the categories described thus far, several broader notions of memorization involve reproducing or inferring general content from the training data, rather than focusing on exact strings (verbatim memorization) or specific entities. This can include, for instance, reproducing factual knowledge in different words than originally presented (often called factual memorization), or recalling underlying concepts without necessarily matching the original phrasing (conceptual memorization or knowledge memorization) [12, 33, 34, 35].
While these types of memorization can be beneficial for tasks requiring reasoning or knowledge application, they also raise distinct concerns. For example, when a model memorizes factual knowledge or concepts, the source of that information becomes critical [36, 37]. If the training data contains inaccuracies, biases, or falsehoods, these errors might be presented in the model outputs as well, leading to trustworthiness issues. Additionally, facts or knowledge could be outdated, further compounding the problem. This is particularly problematic in applications where factual accuracy is paramount, such as educational tools or decision-making systems. Bender et al. [38] highlight how language models can sustain misinformation when trained on unverified or noisy datasets, emphasizing the risks tied to the provenance of memorized knowledge.
II-B Retrievability
@@
A common way to analyze memorization in LLMs is through sampling. This involves selecting tokens probabilistically from the model’s predicted distribution at each step, to see if a sequence of probabilistically generated tokens can be traced back to the training data. Since the LLM’s input and output domains are both discrete, it is important to note that it might not be possible to extract all of the memorized content, given a finite number of generation trials through sampling. Based on how an LLM can be prompted to output the training data, memorization can be classified into extractable and discoverable in terms of retrievability.
@@
Extractable memorization refers to the ability to retrieve specific information from a model’s training data without direct access to that data. Carlini et al. [39] defines extractable memorization as follows:
@@
(Extractable Memorization) Let $LM(p)$ denote the output of a language model when prompted with $p$. An example $s$ from the training set $S$ is extractably memorized if an adversary (without access to $S$) can construct a prompt $p$ that makes the model produce $s$ (that is, $LM(p)=s$).
@@
Discoverable memorization measures the extent to which models can reproduce their training data when explicitly prompted with data from their training set. Nasr et al. [40] suggests the following definition for discoverable memorization:
@@
(Discoverable memorization) Let $LM(p)$ denote the output of a language model when prompted with $p$. For a context $c=(p,s)$ from the training set $S$, we say that $s$ is discoverably memorized if $LM(p)=s$.
A more specific form of this is k-discoverable memorization [^6] Biderman et al. [42] mention this as K-extractable, however, to avoid confusion with the definition of extractability (Definition Definition) we opt to use this terminology for this definition. which adds a criterion to the number of prefix tokens:
Definition.
(k-discoverable memorization) For an example $c=(p,s)$ from the training set, string $s$ is said to be k-discoverable if $s$ is discoverable and $p$ is consisted of $k$ tokens [42].
Considering the limited knowledge of the adversary in the extractable memorization definition, Nasr et al. [40] assume that discoverable memorization provides an upper bound for data extraction. Ideally, discoverable memorization requires querying the model with all of the possible substrings from its entire training set, which is computationally intractable. Also, noteworthy is that discoverable memorization differs from extractable memorization in that the prompt $p$ is known to be from the training set.
@@
Discoverable and extractable Memorization
@@
Remark.
Based on the definitions presented here, one might perceive extractable/discoverable memorization and other concepts within the granularity dimension (section II-A) as equivalent. However, it is important to distinguish them. Extractable and discoverable memorization emphasize the methods used to retrieve memorized samples. In the case of granularity of memorization, the retrieval method is irrelevant; the primary concern is what granularity of the data has been memorized, irrespective of how it is produced by the model. Thus, each of the levels of the granularity, can be encompass both extractable and discoverable memorization.
II-C Desirability
Although memorization of factual information can be helpful for the model to perform more accurately on benchmark tasks such as question answering, other data might be retained without any clear purpose, and this might cause issues with privacy and copyright. In this regard, memorization could be categorized into desirable and undesirable subcategories. [45, 39, 46].
Undesirable memorization
Memorization has been demonstrated to be partly an undesirable phenomenon, often resulting in a range of issues, including:
  • Security vulnerabilities: Unintended memorization of confidential information like API keys or passwords could lead to security breaches [47, 48, 49].
  • Copyright violation: Memorization of copyrighted material may lead to legal challenges, especially if the model can reproduce substantial portions of protected works [16, 50, 51].
  • Bias and fairness: Based on the distribution of the data, memorization could introduce bias issues in the model output.
  • Transparency and auditing: Memorization poses challenges for transparency and auditing by making it difficult to decide whether specific outputs stem from generalization or memorized content, leading to accountability and interpretability issues in LLM systems.
  • Overperformance on the benchmarks: LLMs often memorize benchmark datasets, leading to overfitting and inflated performance on seen data [52, 53].
Desirable memorization
Ranaldi et al. [54] in their experiments show that memorization is could be beneficial for model performance. Moreover, as discussed above, even though unintended, memorization sometimes allows models to store and utilize vast amounts of knowledge, acting as a desirable phenomenon. Key aspects of desirable memorization include:
  • Knowledge retention: LLMs are often deliberately designed to memorize facts, concepts, and general knowledge to enhance their performance in tasks like question answering and information retrieval [55, 56].
Figure 3: The spectrum of memorization could be viewed as a 3-dimensional cube.
Answer to RQ1. By reviewing the literature, we can define memorization across three dimensions: granularity, retrievability, and desirability. These dimensions are illustrated in Figure 3.